# Enhancing reliability in prediction intervals using point forecasters: Heteroscedastic Quantile Regression and Width-Adaptive Conformal Inference

### v1

@@ 4 Real data examples
To test the effectiveness of the final proposed method, WACI-HQR, we will compare the different interval construction schemes presented during the paper. As quantile regression methods, the QRA model and our HQR model will be compared. Also, we will compare the different stages of conformal post-processing on the two quantile regressions considered: the conformalization of the prediction bands, CQR, and the adaptive versions of the same ACI-CQR and WACI-CQR, which will be denoted as ACI and WACI, respectively. In this way, it will be possible to assess the value of each modelling step on the desired properties of the intervals: from the initial quantile regression to the final adaptive conformal model. For the application of WACI the weights shown in (4) are used. The results with the weights described in (5) are almost equivalent.
4.1 Data
4.1.1 Electricity Price Forecasting (EPF)
Data for the Day-Ahead market price in Spain and four different one day-ahead point forecasters ($M=4$, Figure 5) is available from 1st of January 2022 to 5th of October 2023. The period from 5th of October 2022 to 5th of October 2023 is considered as test data. This is the window where conformal methods are applied. In order for these methods to be applied, quantile regression forecasts must be available. These will be obtained from 8th of April 2022 (Figure 6). We believe that this period is ideal for testing this type of models because moments of great uncertainty can be observed at the same time as very steady phases.
Figure 5: One week example of the 4 different point forecasters for the EPF example.
Figure 6: Time series of the Spanish Day-Ahead market between 1st of January 2022 and 5th of October 2023. The predictions of the quantile regression algorithms are obtained from 8th of April 2022 until the end. The Conformal Prediction methods are applied from 5th October 2022.
When forecasting the price one day in advance, it must be predicted for the 24 hours of the next day. Thus, there are two ways of proceeding with the quantile regression: consider 24 daily series (one per hour) and a quantile regression model for each one of these series or a single hourly frequency series and therefore a single quantile regression model for all hours. In any case, a training rolling window of 180 days is considered. That is, at the time of predicting the day $D$, data from days $D-180$ to $D-1$ are considered to train the quantile regression models. For the day $D+1$, the window from $D-179$ to $D$ is considered for training and so on. For those days for which the training window is shorter than 180 days due to data unavailability, all available data will be used. The size of 180 days has been selected empirically after observing the results with several windows. Conformalization is always carried out individually for each hour. For the adaptive methods we will take $\gamma=0.02$, which seems reasonable in view of previous studies (Zaffran et al., 2022) and for the WACI approach $\sigma=3$, which also provides an appropriate performance given the price scale. With all these distinctions, 16 possible methodologies will be compared over a test period of one year. Four possible values of $\alpha$ will be distinguished: $0.2,0.1,0.05,0.01$, which correspond to coverage values of 80, 90, 95 and 99%, respectively.
4.1.2 Wind Power Forecasting (WPF)
Data on wind power generated in Spain and three different 24 hour-ahead point forecasters ($M=3$, Figure 7) is available from 13th of March 2020 to 31st of December 2021. The period from 1st of January 2021 to 31st of December 2021 is considered as test data. Quantile regression forecasts will be obtained from 5th of July 2020 (Figure 8).
Figure 7: One week example of the 3 different point forecasters for the WPF example.
Figure 8: Wind power generation time series between 13th of March 2020 and 31st of December 2023. The predictions of the quantile regression algorithms are obtained from 5th of April 2020 until the end. The Conformal Prediction methods are applied from 1st January 2021.
Unlike the approach for pricing, in the case of wind power production such differentiation between hours is not applied. A global model that takes into account the 24 hours of the day is considered. In this case, where hours are not differentiated, we employ two quantile regression models combined with three conformal post-processing techniques, resulting in a total of eight methodologies to be evaluated. For this example, a training window of 180 days is considered, although notably effective results have also been observed with larger windows.
For those days for which the training window is shorter than 180 days due to data unavailability, all available data will be used. It should be noted that the generation of the last 24 hours is not accessible. For the adaptive methods we will take $\gamma=0.02$ and for the WACI approach $\sigma=0.5$, again due to the scale of the data. Four possible values of $\alpha$ will be distinguished: $0.2,0.1,0.05,0.01$.
@@ 4.2 Evaluation metrics
The empirical coverage is used to measure the validity property: if there are $N$ predictions, the empirical coverage on those predictions is defined as
$\dfrac{1}{N}\sum_{i=1}^{N}\mathds{1}\left\{\hat{l}_{\alpha,i}\leq y_{i}\leq\hat{u}_{\alpha,i}\right\}.$
For an objective miscoverage rate of $\alpha$, the empirical coverage is sought to be as close to $1-\alpha$ as possible. There is a consensus that overcoverage is preferred to undercoverage, so the intervals are valid. The efficiency is usually measured through the mean or median interval length. The adaptive conformal procedures are not limited in the modification of $\alpha$ over time. Although theoretically this makes sense, in practice getting infinite or empty intervals does not add much value. Thus, a practical modification is made to these methods. If the $\alpha$ value of the next iteration is greater than one or less than zero, $\alpha$ is not modified. Thus, the mean and median of the interval length are also used. These efficiency measures, while informative, do not allow comparison of different methodologies unless they produce the same level of empirical coverage, which is unlikely. The Winkler score (Winkler, 1972) is used to measure validity and efficiency together. For each time step $t$ and for a miscoverage rate of $\alpha$, it is defined as the length of the interval plus a penalty term proportional to how far the prediction is from being in the interval:
$W_{\alpha,t}=\begin{cases}(\hat{u}_{\alpha,t}-\hat{l}_{\alpha,t})+\frac{2}{\alpha}(\hat{l}_{\alpha,t}-y_{t})\quad&\text{ if }y_{t}<\hat{l}_{\alpha,t}\\ (\hat{u}_{\alpha,t}-\hat{l}_{\alpha,t})\quad&\text{ if }\hat{l}_{\alpha,t}\leq y_{t}\leq\hat{u}_{\alpha,t}\\ (\hat{u}_{\alpha,t}-\hat{l}_{\alpha,t})+\frac{2}{\alpha}(y_{t}-\hat{u}_{\alpha,t})\quad&\text{ if }y_{t}>\hat{u}_{\alpha,t}\\ \end{cases}$
@@
The main limitation of the Winkler score is that it lacks intuitive interpretability when considered in isolation; it primarily serves as a comparative metric for assessing different prediction intervals. While the mean and median lengths of the intervals can offer some insight into their behaviour, these statistics alone do not provide a comprehensive understanding of how the intervals function. The quantile (X-axis) vs. inverse quantile of the interval length (Y-axis) plot will also be performed to observe the distribution of the interval length and to observe its variance and behaviour beyond the mean/median situation. This chart is in fact a reduced version of the inverse empirical interval length distribution function for each method that allows for a better comparison of the methods with each other.
Getting intervals that vary depending on the error of the point forecasting model is an important characteristic. Thus, groups of observations of the same size will be made, having previously ordered them by interval length [^4] The quantiles of the interval length will be used to create these groups. That is, the observations with interval size 0% to the largest 5%, from 5% to 10%, from 10% to 15% and so on.. The mean absolute error (MAE) for point predictions on each of these groups will be computed. If the error is increasing between groups, then the intervals are informative and are a first step to be able to use them correctly in decision support.
Similarly, to assess if coverage is independent of the length of the interval, these groups will be made again and the coverage in each of them will be checked. The mean deviation of each of these groups from the target coverage and the maximum deviation will be calculated as metrics to measure this property. As these measures depend on the formation of the groups, different group sizes will be taken into account: groups consisting of 5% and 10% of the observations will be considered.
Since in the EPF example 24 different forecasts are being evaluated at each step, the hourly empirical coverage will also be measured for this particular case.
5 Results and discussion

### v2

@@ 4 Computational experiments
To evaluate the effectiveness of the proposed WACI-HQR method, the different interval construction schemes are compared. The quantile regression models include the QRA model, the HQR model, and the HQR-W model. Additionally, the conformal post-processing methods, ACI and WACI, are applied to these models. This comparison evaluates the impact of each modelling step on the interval properties, from the initial quantile regression to the final adaptive conformal approach. The weights used for WACI are given in equation (7), with similar results observed using the weights in equation (8).
@@ 4.1 Evaluation metrics
Mean empirical coverage
The mean empirical coverage is used to measure the validity property: if there are $N$ prediction intervals for observations $y_{i},\,i=1,\dots,n$, the empirical coverage on those predictions is defined as
$\frac{1}{N}\sum_{i=1}^{N}\mathds{1}\left(y_{i}\in\widehat{C}_{\alpha,i}\right).$
For an objective miscoverage rate of $\alpha$, the empirical coverage is sought to be as close to $1-\alpha$ as possible.
Mean interval length
The efficiency is usually measured through the mean or median interval length. The ACI and WACI adaptive conformal procedures are not constrained in their definition of $\alpha$, allowing for intervals that are either empty sets or cover all of $\mathbb{R}$. While empty intervals are rare, occurring only when the $\alpha$ value in the corresponding iteration exceeds 1, the case of infinite intervals is more frequent, arising when $\alpha$ is less than 00. To address these situations, it is common practice to use the median interval length, as it is unaffected by infinite intervals. In this work, we compute the mean interval length, replacing infinite intervals with a fixed interval defined by the largest upper bound and smallest lower bound of the base model observed in the training set. This approach effectively penalizes the production of excessive infinite intervals.
Winkler score
The Winkler score (Winkler, 1972) is used to measure validity and efficiency together. For each time step $t$ and for a miscoverage rate of $\alpha$, it is defined as the length of the interval plus a penalty term proportional to how far the prediction is from being in the interval:
$W_{\alpha,t}=\begin{cases}(\widehat{u}_{\alpha,t}-\widehat{l}_{\alpha,t})+\frac{2}{\alpha}(\widehat{l}_{\alpha,t}-y_{t})\quad&\text{ if }y_{t}<\widehat{l}_{\alpha,t}\\ (\widehat{u}_{\alpha,t}-\widehat{l}_{\alpha,t})\quad&\text{ if }\widehat{l}_{\alpha,t}\leq y_{t}\leq\widehat{u}_{\alpha,t}\\ (\widehat{u}_{\alpha,t}-\widehat{l}_{\alpha,t})+\frac{2}{\alpha}(y_{t}-\widehat{u}_{\alpha,t})\quad&\text{ if }y_{t}>\widehat{u}_{\alpha,t}\\ \end{cases}$
@@
Pearson’s correlation
Following Feldman et al. (2021), conditional coverage is assessed by computing Pearson’s correlation coefficient between the interval length and the coverage indicator function. The closer to 0, the better, as the indicator function of coverage should be independent of the length when true quantiles are considered.
ILS $\lambda$ Coverage
A variation of the metric $\Delta$ILS-Coverage presented in Feldman et al. (2021) is used to check the effectiveness of the post-processing methodologies. To evaluate whether the modifications made to the intervals produced by quantile regression algorithms are truly useful, consider a base quantile regression model with produces a prediction interval $\widehat{C}_{i}$ and a conformal procedure applied post hoc to this base model which produces the prediction interval $\widehat{C}_{i}^{c}$ for observation $i,\,i=1,\dots,N$. Let $\Delta_{i}$ denote the difference in interval lengths proposed by the two algorithms for observation $i$, defined as
$\Delta_{i}=\left\lvert\lvert\widehat{C}^{c}_{i}\rvert-\lvert\widehat{C}_{i}\rvert\right\rvert.$
Now consider the $\lambda\cdot 100\%$ of samples most affected by the conformal procedure, i.e.,
$\text{ILS}=\left\{i\,:\,\Delta_{i}\geq q_{\lambda}\left(\left\{\Delta_{i}\right\}_{i=1}^{N}\right)\right\},$
where $q_{\lambda}$ denotes the $\lambda$-empirical quantile. The ILS $\lambda$ Coverage metric is then defined as
ILS $\text{ILS }\lambda\text{ Coverage}=\frac{1}{\left|\text{ILS}\right|}\sum_{i\in\text{ILS}}\left|\mathds{1}\left(y_{i}\in\widehat{C}^{c}_{i}\right)-\left(1-\alpha\right)\right|.$
The objective is to determine whether the intervals that have been modified to a greater extent indeed achieve the desired level of coverage. [^2] The difference with Feldman et al. (2021) is that it evaluates only the intervals that have increased in size while ignoring those that have been reduced by the conformal procedure. We propose considering both types of modifications, since reducing the size of the interval when it does not apply should also be penalized. We consider $\lambda=0.10$.
Spearman’s correlation
The Spearman’s correlation coefficient between the mean absolute error of the mean prediction of the point forecasts and the interval length will be computed to assess the strength of the relationship between prediction difficulty and interval length. Since this relationship does not need to be linear, Spearman’s rank correlation coefficient is preferred over the traditional linear correlation coefficient.
Standard deviation of the interval length
The standard deviation of the interval lengths generated by an algorithm is computed to assess the algorithm’s ability to distinguish between varying uncertainty contexts. This metric is closely related to the previous one, as different levels of prediction difficulty are associated with corresponding variations in the interval lengths.
MCD $\lambda$
Let $1-\alpha$ denote the target coverage level, and let’s assume there are $N$ observations for which a prediction interval $\left\{\widehat{C}_{\alpha,i}\right\}_{i=1}^{N}$ has been built. To assess the property described in equation (2), we define the mean coverage deviation (MCD) as follows: the data is divided into $K=\frac{100}{\lambda}$ subsets $\left\{\mathcal{G}_{k}\right\}_{k=1}^{K}$, where each subset contains approximately $\lambda$of the data. The subsets are defined by the empirical quantiles of $\left\{\left|\widehat{C}_{\alpha,i}\right|\right\}_{i=1}^{N}$, such that:
$\mathcal{G}_{k}=\left\{i\,:\,q_{\frac{k-1}{K}}\left(\left|\widehat{C}_{\alpha,i}\right|\ \right)\leq\left|\widehat{C}_{\alpha,i}\right|\ <q_{\frac{k}{K}}\left(\left|\widehat{C}_{\alpha,i}\right|\ \right)\right\},\quad k=1,\dots,K.$
For each subset $\mathcal{G}_{k}$, the deviation between its mean empirical coverage and the objective coverage is computed
$D_{k}=\left|\frac{1}{\left|\mathcal{G}_{k}\right|}\sum_{i\in\mathcal{G}_{k}}\mathds{1}\left(y_{i}\in\widehat{C}_{\alpha,i}\right)-\left(1-\alpha\right)\right|.$
Finally, the mean coverage deviation is defined as
$\text{MCD}=\frac{1}{K}\sum_{k=1}^{K}D_{k}.$
A smaller value of MCD indicates better alignment of empirical coverage with the target coverage, regardless of interval length. We consider $\lambda=5$.
We propose two metrics that, to the best of our knowledge, have not been utilized in similar studies: Spearman’s correlation coefficient and the MCD. These metrics provide valuable insights into the performance of the various methods.
4.2 A synthetic example
A synthetic example is designed to evaluate the effectiveness of WACI and provide insights into its behaviour. Consider a time series $y_{t}$ generated from a normal distribution $N(\mu,\sigma_{t})$, where the standard deviation $\sigma_{t}$ alternates between two states representing different levels of uncertainty. Specifically, the process alternates between:
  • 1.
    A high-uncertainty state where $\sigma_{t}=\sigma_{1}=7$,
  • 2.
    A low-uncertainty state where $\sigma_{t}=\sigma_{2}=2$.
The transition between these two states is governed by a probabilistic mechanism. The process begins in the high-uncertainty state ($\sigma_{t}=\sigma_{1}$), and at each time step, the probability of transitioning to the other state increases incrementally by 0.0001. Once a transition occurs, the probability resets to zero, ensuring alternating states. A binary indicator variable $\delta_{t}$ is used to represent the current state: $\delta_{t}=0$ corresponds to the high-uncertainty state, while $\delta_{t}=1$ represents the low-uncertainty state.
For simplicity, the mean $\mu=100$ is assumed to be known. To compute the length of the interval it is simulated that a sample of size 10 is drawn from the distribution at each time step $t$. However, instead of estimating the standard deviation from a sample, a deterministic relationship is used for $\widehat{\sigma}_{t}$, introducing smooth time-dependent fluctuations that reflect transitions between overcoverage and undercoverage. This is given by:
$\widehat{\sigma}_{t}=\begin{cases}\sigma_{1}+2\cdot\sin(0.001\cdot t)&\text{if}\delta_{t}=0,\\ \sigma_{2}+\cos(0.005\cdot t)&\text{if }\delta_{t}=1.\end{cases}$
and the unconformalized interval is computed based on the following relationship:
$\widehat{l}_{\alpha,t}=\mu-T_{1-\frac{\alpha}{2},9}\cdot\widehat{\sigma_{t}}\cdot\sqrt{1+\frac{1}{10}},\quad\widehat{u}_{\alpha,t}=\mu+T_{1-\frac{\alpha}{2},9}\cdot\widehat{\sigma_{t}}\cdot\sqrt{1+\frac{1}{10}},$
where $T_{1-\frac{\alpha}{2},9}$ is the critical value of the $t$-distribution for a two-tailed test with significance level $\alpha$ and 9 degrees of freedom.
This design ensures smooth changes in coverage, making the example particularly well-suited to evaluating ACI-based conformalization methods.
Figure 6 illustrates 1000 time steps of this process, showing the generated observations, true intervals, and unconformalized intervals.
Figure 6: Simulated time series data, true intervals, and unconformalized intervals over 1000 time steps.
The goal is to conformalize the intervals such that they remain valid and efficient across both uncertainty states, while achieving optimal performance according to the evaluation metrics introduced in Section 4.1. For this experiment, hyperparameters were fixed ($\gamma=0.01$ for ACI and WACI, $\sigma=1$ for WACI), and no optimization was performed. To account for the randomness inherent in the process, the experiment was repeated 100 times with different seeds, enabling standard deviation estimates for all metrics. Each run simulated a time series of length 10000 with $\alpha=0.2$.
4.2.1 Results and discussion
Tables 1 and 2 present the values of the different evaluation metrics previously described for the two uncertainty states separately.
Method
Mean Empirical
Coverage
Average
Length
Winkler
Score
Pearson
Correlation
ILS 0.10 MCD 5
Initial 85.13 (0.68) 21.14 (0.23) 26.05 (0.22) 0.25 (0.01) 9.17 (0.52)
ACI 83.24 (0.51) 19.90 (0.17) 25.48 (0.24) 0.22 (0.01) 3.30 (0.55) 7.25 (0.50)
WACI 81.08 (0.23) 18.59 (0.23) 24.89 (0.27) 0.15 (0.01) 1.47 (0.30) 4.35 (0.37)
Table 1: Mean results of 100 runs of the synthetic experiment for the high uncertainty state samples. The standard deviation of the metrics is shown in brackets.
Method
Mean Empirical
Coverage
Average
Length
Winkler
Score
Pearson
Correlation
ILS 0.10 MCD 5
Initial 79.92 (1.19) 5.88 (0.12) 7.97 (0.10) 0.34 (0.02) 14.27 (0.83)
ACI 76.64 (0.59) 5.13 (0.09) 7.57 (0.10) 0.25 (0.02) 3.34 (0.70) 9.87 (0.59)
WACI 80.72 (0.13) 5.35 (0.07) 7.18 (0.09) 0.10 (0.01) 0.92 (0.22) 4.57 (0.36)
Table 2: Mean results of 100 runs of the synthetic experiment for the low uncertainty state samples. The standard deviation of the metrics is shown in brackets.
A quick comparison of the classical empirical coverage and mean interval length values reveals that WACI is the only method that effectively captures both states. This observation is further supported by the Winkler Score, which is the lowest in both cases. The other metrics provide deeper insights into why WACI is clearly the preferred choice. Notably, WACI does not exhibit the same dependence between interval length and coverage as the other two intervals, as shown in the values of the Pearson’s correlation coefficient. Specifically, the large changes observed in ACI do not align closely with the desired coverage levels, unlike WACI. This difference is evident in the ILS 0.10 metric. The MCD metric highlights how WACI achieves the desired outcome by reducing the dependence between interval length and empirical coverage, resulting in intervals with superior overall characteristics. In particular, WACI allows for targeted adjustments at each uncertainty state without the need for incremental fine-tuning, as seen with ACI. This distinction is more clearly illustrated in Figure 7.
Figure 7: Prediction intervals produced by ACI (green) and WACI (orange) over the synthetic example.
Figure 7 illustrates how, following a change in behaviour, ACI requires several iterations to detect and adjust to the necessary modifications for approximating the true interval. This behaviour is expected since ACI relies solely on information from the previous iteration to guide its corrections. In contrast, WACI adapts immediately to the change, as it is capable of distinguishing between the two behavioural states present in the system. This is because WACI references prior iterations where the initial interval length, before conformalization, was similar.
It is essential to analyse the two (or more) states separately when their existence is known. Table 3 presents the results of the evaluation metrics when all observations are considered together. At first glance, ACI appears to perform well, achieving results only slightly inferior to WACI in terms of mean empirical coverage and mean interval length. However, as shown in Tables 1 and 2, this overall performance obscures significant discrepancies: in some situations, ACI exhibits over-coverage, while in others, there is clear under-coverage, leading to the observed average outcomes. Additionally, the other metrics reinforce that these discrepancies are closely tied to the relationship between interval length and coverage. The only metric for which ACI performs well is ILS 0.10, but this result warrants closer examination, particularly to identify which observations are most significantly modified by ACI.
Method
Mean Empirical
Coverage
Average
Length
Winkler
Score
Pearson
Correlation
ILS 0.10 MCD 5
Initial 82.51 (0.75) 13.48 (0.44) 19.68 (0.52) 0.17 (0.02) 11.13 (0.55)
ACI 79.93 (0.02) 12.49 (0.41) 16.49 (0.52) 0.15 (0.01) 0.09 (0.07) 7.89 (0.42)
WACI 80.90 (0.11) 11.95 (0.40) 16.01 (0.52) 0.04 (0.003) 1.17 (0.17) 3.68 (0.29)
Table 3: Mean results of 100 runs of the synthetic experiment for every observation. The standard deviation of the metrics is shown in brackets.